OpenStudy (anonymous):

how do you solve this proportion 6/9 = x+4/12

OpenStudy (anonymous):

First, multiply both sides by 36.

OpenStudy (anonymous):

i dont think thats how you do it

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

6/9 = 2/3

OpenStudy (anonymous):

nv

OpenStudy (anonymous):

is it (x+4)/12, or x + (4/12)?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

4/12 = 1/3

OpenStudy (anonymous):

you have to simplify first

OpenStudy (anonymous):

and then left the x alone

OpenStudy (anonymous):

2/3 - 1/3 = 1/3 = x

OpenStudy (anonymous):

So, you have: (x+4)/12 = 6/9 (x+4) = 6 * 12 / 9 x + 4 = 8 x = 8-4 x = 4

OpenStudy (anonymous):

if it is with the parentesis, it is different. 2/3 = (x+4)/12 you simplify by 3 both sides, and multiply by 4, so 2*4 = x+4, and X = 4

OpenStudy (anonymous):

you have 6/9=(x+4)/12 separate what's on the right of the equal sign into 2 fractions; x/12 +4/12 subtract 4/12 from both sides and you get 1/3 = x/12 you then multiply both sides by 12 and get 4=x is that your answer?